L passes through (0,-3) and is parallet to L2: 4x-3y+0=0. L has equation y=mx+b, where m=? b=?
Answer by mananth(16949)

4 x -3 y = 0
Find the slope of this line
make y the subject
-3 y = -4 x + 0
Divide by -3
y = 4/3x
compare with y=mx+b
slope m = 4/3
The slope of a line parallel to the above line will be the same
m= 4/3 ,point ( 0 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = 0 + b
b= -3
m= 4/3
Plug value of the slope and b in y = mx +b
The required equation is y = 4/3x-3

m.ananth@hotmail.ca